# SATA-Bench: Select All That Apply Benchmark for Multiple Choice Questions

### v2

@@ H.3.1 Few Shot prompt
[…]
@@ H.4 Think Option by Option prompt
Inspired by [46, 39], we instruct LLM to understand each options and analyze each answer independently.
[…]
@@ H.4.1 Few Shot Option prompt
[…]
@@ H.4.2 Prompt with Average Options Count
[…]
@@ H.4.3 Prompt with Correct Number of Options
[…]
@@ H.4.4 Single Choice Prompt
[…]
@@ H.5 Prompt with Numeric Option
[…]
@@ H.6 Prompt with small alphabet Option
[…]
@@ Appendix I Inference Error Handling
The original PriDe algorithm [56] is designed for processing MCQ question sets with fixed option set length (usually 4). It works by observing the probability changes when performing permutations of option IDs for each question, and it can compute priors, which is known as the probabilistic mass that the model a priori assigns to option ID tokens.
@@
Option “B" becomes top-1 after we remove the heavy prior on “A". To learn more low-level details, please refer to the original paper [56]).
@@
We chose a fixed 90% confidence threshold as the stopping condition (ii) in Choice Funnel for all models. While this initial parameter selection was chosen for its simplicity, later evaluation indicated that it yielded sufficiently robust performance. Consequently, we did not pursue further investigation into more granular threshold adjustments. It also demonstrates that the algorithm is generalizable to other models without careful calibration.
The first baseline method first token sets a fixed threshold so that any option with a probability above the threshold is selected, and this should be the lower bound of the performance. First token debiasing can be used to find out if the popular strategy used to solve the MCQ questions is transferable to the SATA questions in terms of minimizing the impact of the selection bias. Lastly, we expect yes/no to be a competitive baseline given that it processes each choice separately.
@@
Models. Our study focuses on the causal, decoder-only LLMs since this architecture has become the dominant choice for modern LLMs. We experiment with 7 LLMs from Table 2 under Probability Based Retrieving which are all popular open-source models on the HuggingFace website, and we can access their output probabilities: DeepSeek R1 Distilled LLAMA 8B [14], Qwen2.5 14B [54], Ministral 8B [50], Phi 3 7B [1], Phi 4 mini reasoning [2], Bloomz 7B [35], and Llama 3.1 8B [52].
@@
We compared two commonly employed auxiliary response options in traditional survey science domain [45]: ’I don’t know’ (IDK) and ’None of the above’ (NOTA), examining their effectiveness as Choice Funnel stopping condition. Based on an ablation study on Table 13, NOTA yields consistently better performance. When using IDK, we observe noticeable increase in InfCost and result in worse Count Bias (CtDifAbs and CtAcc), which means model tends to over select number of options, indicating that the model would rather select a wrong answer than saying “I don’t know". This is potentially related to RLHF process, where the model is trained to generate answers that are more favorable to humans.
@@
We conducted an ablation study on the two sub-components of Choice Funnel: token debiasing ("debiasing only") and iterative selection (the process of iteratively selecting options until a stopping condition is met, denoted as "CF only"). The analysis is performed on 3 open-source models. When comparing "CF only" to the complete "CF + debiasing", the observed increase in SPD metric demonstrates that token debiasing effectively mitigates unselection bias, yielding better performance. Nevertheless, the comparison between "debiasing only" and "CF only" reveals that our novel iterative selection component contributes more substantially to overall performance improvements.
@@
All hyper-parameters, prompts, and decoding settings were kept identical to the main benchmark; only the answer order was shuffled once per question. Table 14 reports the difference (permuteoriginal) for each metric, so negative values indicate a drop in performance and positive values indicate an increase. CtDif is shown with a downward arrow even though its baseline values are negative; a more negative CtDif therefore indicates a larger absolute mismatch in option counts.
@@
Claude 3 Haiku $-24.06$ $-34.69$ $-34.28$ $-35.31$ $+6.06$ $+0.17$ $+0.12$ $-0.07$ $-0.51$
Llama 3.1 405B $-3.80$ $-3.90$ $-4.71$ $-5.22$ $+9.73$ $-0.20$ $+0.25$ $-0.18$ $-0.71$
@@
All three models suffer performance degradation when answer choices are shuffled, with Claude 3 Haiku exhibiting the sharpest decline (–24 EM, –35 JI). Selection / count-bias metrics (RStd, SPD, CtDifAbs) increase for every model except RSD, confirming heightened positional bias.
@@ Discussion.
These results suggest that current LLMs implicitly learn positional heuristics from training data in which answer orders are fixed. Breaking this assumption makes the models less certain and more prone to biased guessing. Future work should examine (i) whether fine-tuning on randomly ordered choices mitigates the effect, and (ii) how pronounced the bias is for other model families and task domains.

### v3

@@ H.3.1 Few Shot prompt
Few Shots Prompt Given the following question and four candidate answers (A, B, C, and D), choose the best answer. Question 1: question 1 Option 1: option 1 Answer 1:correct option json 1 Question 2: question 2 Option 2: option 2 Answer 2: corect option json2 Question 5: question 5 Option 5: option 5 Answer 5:correct option json 5 Question: question Option: option Please select all choices that apply. You must focus on the question and select all choices that apply. Let’s think step by step: You must present your selected option IDs in the following JSON format: $\{{"choices":<A|B|C|D|E|F|G|H|I|J|K|L|M|N|O>}\}$
@@ H.4 Think Option by Option prompt
Inspired by Smyth et al. (2006); Pew Research Center (2019), we instruct LLM to understand each options and analyze each answer independently.
Choice-by-choice Prompt Given the following question and four candidate answers (A, B, C, and D), choose the best answer. Question: question Option: option Let’s think through this step by step: 1. First, let’s understand what the question is asking… 2. Now, let’s evaluate each option individually… 3. Therefore, the correct answers are… You must present your selected option IDs in the following JSON format: $\{{"choices":<A|B|C|D|E|F|G|H|I|J|K|L|M|N|O>}\}$
@@ H.4.1 Few Shot Option prompt
Few Shots Option Prompt Given the following question and four candidate answers (A, B, C, and D), choose the best answer. Question 1: question 1 Option 1: option 1 Choice by choice reasoning 1: reason 1 Answer 1:correct option json 1 Question 2: question 2 Option 2: option 2 Choice by choice reasoning 2: reason 2 Answer 2: corect option json2 Question 5: question 5 Option 5: option 5 Choice by choice reasoning 5: reason 5 Answer 5:correct option json 5 Question: question Option: option Let’s think through this step by step: 1. First, let’s understand what the question is asking… 2. Now, let’s evaluate each option individually… 3. Therefore, the correct answers are… You must present your selected option IDs in the following JSON format: $\{{"choices":<A|B|C|D|E|F|G|H|I|J|K|L|M|N|O>}\}$
@@ H.4.2 Prompt with Average Options Count
SATA Prompt Given the following question where there is more than one correct answer, choose all correct answers. Question: question Choices: choices Please select all choices that apply. You must focus on the question and select all choices that apply. The number of average selected options is $3.63$. Let’s think step by step: You must present your selected option IDs in the following JSON format: $\{{"choices":<A|B|C|D|E|F|G|H|I|J|K|L|M|N|O>}\}$
@@ H.4.3 Prompt with Correct Number of Options
SATA Prompt Given the following question where there is more than one correct answer, choose all correct answers. Question: question Choices: choices Please select all choices that apply. You must focus on the question and select all choices that apply. The number of average selected options is $X$. Let’s think step by step: You must present your selected option IDs in the following JSON format: $\{{"choices":<A|B|C|D|E|F|G|H|I|J|K|L|M|N|O>}\}$
@@ H.4.4 Single Choice Prompt
Single Choice Prompt Given the following question where there is only one correct answers, choose the correct answer. Question: question Choices: choices Please the correct choice that apply. Let’s think step by step: You must present your selected option IDs in the following JSON format: $\{{"choice":<A|B|C|D|E|F|G|H|I|J|K|L|M|N|O>}\}$
@@ H.5 Prompt with Numeric Option
Numeric Prompt Given the following question where there is more than one correct answer, choose all correct answers. Question: question Choices: choices Please select all choices that apply. You must focus on the question and select all choices that apply. You must present your answers in ascending orders. Let’s think step by step: You must present your selected option IDs in the following JSON format: $\{{"choices":<1|2|3|4|5|6|7|8|9|10|11|12|13|14|15>}\}$
@@ H.6 Prompt with small alphabet Option
Small Alphabet Prompt Given the following question where there is more than one correct answer, choose all correct answers. Question: question Choices: choices Please select all choices that apply. You must focus on the question and select all choices that apply. Let’s think step by step: You must present your selected option IDs in the following JSON format: $\{{"choices":<a|b|c|d|e|f|g|h|i|j|k|l|m|n|o>}\}$
@@ Appendix I Inference Error Handling
The original PriDe algorithm (Zheng et al., 2024) is designed for processing MCQ question sets with fixed option set length (usually 4). It works by observing the probability changes when performing permutations of option IDs for each question, and it can compute priors, which is known as the probabilistic mass that the model a priori assigns to option ID tokens.
@@
Option “B" becomes top-1 after we remove the heavy prior on “A". To learn more low-level details, please refer to the original paper (Zheng et al., 2024)).
@@
We chose a fixed 90% confidence threshold as the stopping condition (ii) in Choice Funnel for all models. This initial parameter selection was tuned on 100 hold out data points from raw dataset instead of evaluation set and moves to the closest number that can be divided by 10. It demonstrates that the algorithm is generalizable to other models without careful calibration.
The first baseline method first token sets a fixed threshold so that any option with a probability above the threshold is selected, and this should be the lower bound of the performance. First token debiasing can be used to find out if the popular strategy used to solve the MCQ questions is transferable to the SATA questions in terms of minimizing the impact of the selection bias. Lastly, we expect yes/no to be a competitive baseline given that it processes each choice separately with cost of increased inference compute.
@@
Models. Our study focuses on the causal, decoder-only LLMs since this architecture has become the dominant choice for modern LLMs. We experiment with 7 LLMs from Table 2 under Probability Based Retrieving which are all popular open-source models on the HuggingFace website, and we can access their output probabilities: DeepSeek R1 Distilled LLAMA 8B (DeepSeek-AI et al., 2025), Qwen2.5 14B (Yang et al., 2025), Ministral 8B (Team, 2024), Phi 3 7B (Abdin et al., 2024), Phi 4 mini reasoning (Abdin et al., 2025), Bloomz 7B (Muennighoff et al., 2022b), and Llama 3.1 8B (Touvron et al., 2023).
@@
We compared two commonly employed auxiliary response options in traditional survey science domain (Schuman & Presser, 1996): ’I don’t know’ (IDK) and ’None of the above’ (NOTA), examining their effectiveness as Choice Funnel stopping condition. Based on an ablation study on Table 14, NOTA yields consistently better performance. When using IDK, we observe noticeable increase in InfCost and result in worse Count Bias (CtDifAbs and CtAcc), which means model tends to over select number of options, indicating that the model would rather select a wrong answer than saying “I don’t know". This is potentially related to RLHF process, where the model is trained to generate answers that are more favorable to humans.
@@
The CF only setting represents scenarios where the model has no access to raw probabilities and instead relies solely on the Choice Funnel algorithm (Black-box settings). Compared to token debiasing, this approach achieves significant improvements in EM and Precision. On average, across three models—even without using token probabilities —Choice Funnel yields a 10.79% increase in Exact Match, a 20.51% increase in Jaccard Index, a 13.4 reduction in SPD, and a 0.86 reduction in CtAbsDif.
We conducted an ablation study on the two sub-components of Choice Funnel: token debiasing ("debiasing only") and iterative selection (the process of iteratively selecting options until a stopping condition is met, denoted as "CF only"). The analysis is performed on 3 open-source models.
When comparing "CF only" to the complete "CF + debiasing", the observed increase in SPD metric demonstrates that token debiasing effectively mitigates unselection bias, yielding better performance. Nevertheless, the comparison between "debiasing only" and "CF only" reveals that our novel iterative selection component contributes more substantially to overall performance improvements.
@@
M.4 Scalability of Choice Funnel on Larger LLM
Table 15: Scalability demonstration with larger LLAMA3.1-70B model, showing that ChoiceFunnel improves performance across different model sizes.
[…]
LLAMA3.1-70B + prompting 17.94 60.64 1.81 0.22
LLAMA3.1-7B + ChoiceFunnel 19.88 56.19 7.75 0.33
LLAMA3.1-70B + ChoiceFunnel 24.43 68.66 0.37 0.37
These results show that Choice Funnel scales well with model size, and consistently outperforms prompting-only approaches while maintaining high efficiency.
@@
All hyper‑parameters, prompts, and decoding settings were kept identical to the main benchmark; only the answer order was shuffled once per question. Table 16 reports the difference (permuteoriginal) for each metric, so negative values indicate a drop in performance and positive values indicate an increase. CtDif is shown with a downward arrow even though its baseline values are negative; a more negative CtDif therefore indicates a larger absolute mismatch in option counts.
@@
Claude 3 Haiku 24.06-24.06 34.69-34.69 34.28-34.28 35.31-35.31 +6.06+6.06 +0.17+0.17 +0.12+0.12 0.07-0.07 0.51-0.51
Llama 3.1 405B 3.80-3.80 3.90-3.90 4.71-4.71 5.22-5.22 +9.73+9.73 0.20-0.20 +0.25+0.25 0.18-0.18 0.71-0.71
@@
All three models suffer performance degradation when answer choices are shuffled, with Claude 3 Haiku exhibiting the sharpest decline (–24 EM, –35 JI). Selection / count‑bias metrics (RStd, SPD, CtDifAbs) increase for every model except RSD, confirming heightened positional bias.
@@ Discussion.
These results suggest that current LLMs implicitly learn positional heuristics from training data in which answer orders are fixed. Breaking this assumption makes the models less certain and more prone to biased guessing. Future work should examine (i) whether fine‑tuning on randomly ordered choices mitigates the effect, and (ii) how pronounced the bias is for other model families and task domains.
@@
Appendix P The Challenge of Multi-Answer Reasoning
P.1 Problem Setup
We formalize SATA questions as a subset prediction task. Given a set of $K$ candidate options $\mathcal{O}=\{o_{1},\ldots,o_{K}\}$ and a ground-truth set $S^{\star}\subseteq\mathcal{O}$ of correct options, a model must output $\hat{S}\subseteq\mathcal{O}$ that matches $S^{\star}$. We evaluate with set-based metrics including exact match (EM), Jaccard index (JI), macro precision/recall, and count-based measures (count difference, absolute count difference, and count accuracy; see Appendix F for definitions). Unlike single-choice MCQ (where $|S^{\star}|=1$), SATA requires reasoning over both which options are correct and how many should be selected.
P.2 Bias Definitions
Let $y_{i}^{\star}\in\{0,1\}$ denote the ground-truth label for option $o_{i}$ and $\hat{y}_{i}\in\{0,1\}$ the model’s selection. Define the random variables $C^{\star}=\sum_{i=1}^{K}y_{i}^{\star}$ and $\hat{C}=\sum_{i=1}^{K}\hat{y}_{i}$ as the true and predicted counts.
Count Bias.
A model exhibits count bias if it systematically under estimates the number of correct options: $\mathbb{E}[\hat{C}]\neq\mathbb{E}[C^{\star}]$ over the evaluation distribution. Empirically, we find a dominant under-selection pattern, $\mathbb{E}[\hat{C}]<\mathbb{E}[C^{\star}]$ (Sec. 3.2; Figures 12, 11), reflected in low CtAcc and negative mean CtDif.
Selection Bias.
Let $p_{i}=\Pr(\hat{y}_{i}=1)$ denote the marginal selection probability for option $o_{i}$ across the benchmark. A model exhibits selection bias if the dispersion of $\{p_{i}\}_{i=1}^{K}$ is larger than expected from the true label distribution, indicating preference or aversion to certain labels independent of content. [^3] Position and formatting effects can contribute; cf. Zheng et al. (2024). We quantify selection skew with RStd/RSD (Zheng et al., 2024; Croce et al., 2020; Reif & Schwartz, 2024) and introduce Selection Probability Divergence (SPD) to capture unselection bias (Appendix F); in aggregate, observed SPD significantly exceeds random baselines (Sec. 3.2).
Speculation Bias.
Define the per-question false-positive count $\mathrm{FP}=\sum_{i=1}^{K}(1-y_{i}^{\star})\hat{y}_{i}$ and the speculation indicator $FPR=\mathbb{E}[\mathrm{FP}>0]$. A model exhibits speculation bias if it systematically selects options outside the gold set, especially more than the number of time it produces correct labels, $FPR>EM$. Speculation bias is reflected by higher macro false-positive rate and smaller JI (which penalizes any spurious selections). Note that speculation may co-occur with over-selection, but it is distinct: a model can be count-unbiased yet still speculate (high FPR).
Appendix Q Does LRM help? A case study of GPT-OSS on SATA-Bench
Reasoning model such as GPT-OSS 120B model performs on par with GPT-4.1 on SATA-Bench. GPT-OSS 20B model is much weaker than 120B but still matches Llama-3.1-405B. Despite good slightly better performance. Reasoning model does exhibit a few failure modes in SATA-Bench.
Repetitive Reasoning. We define a reasoning as repetitive if it repeats 100+ characters more than 10 times. This happens in 11% of reasons for 20B model. In those cases where model produce repetitive reasoning, it have much lower EM rate. As an example, GPT-OSS 20B exact match rate drop form 27.4% to 18.5% when it starts to repeat the same characters.
Reason Answer Mismatch. The final answer choices do not always align with the reasoning steps. We used Claude 3 Haiku to extract answers. We found that for cases where 120B is correct and 20B is wrong, 53.2% of the answers do not match the reasoning. In 45% of mismatched cases, the reasoning itself was actually correct — the model just picked the wrong a subset of correct choices as final answer. This has increase unselection bias and count bias. We provide the following as an example of GPT-OSS 120B where there is a mismatch
Example: Reason–Answer Mismatch Case Context. Sunlight Sunlight can be used to heat homes. It can also be used to produce electricity. This conversion is made possible by solar cells. However, solar energy may not always be practical. Some areas are just too cloudy. Example Solar panels on the roof of this house generate enough electricity to supply a family’s needs. Moving Water Falling water can have a lot of energy. Its energy can be converted into kinetic energy. This energy can turn a turbine and generate electricity. The water may fall naturally over a waterfall or flow through a dam. A drawback of dams is that they flood land upstream. They can also reduce water flow downstream. Either effect may harm ecosystems. Wind is moving air. It has kinetic energy that can do work. Wind turbines change the kinetic energy of the wind to electrical energy. Only certain areas of the world get enough steady wind. Many people also think that wind turbines are noisy and not very nice to look at. Question. How does moving water make electricity? (Select all that apply) Options. A. Water can generate electricity by turning a turbine. B. Electricity can be made by storing the water in batteries. C. Water’s energy can turn a turbine and magically result in electricity. D. Falling water can have a lot of energy which can be converted into kinetic energy that turns a turbine and generates electricity. E. Water can move a turbine that generates power. F. Electricity can be made by taking the hydrogen out of moving water. Model Reasoning (excerpt). Focus on the “Moving Water” sentences: (7) Falling water can have a lot of energy; (8) its energy can be converted into kinetic energy; (9) this energy turns a turbine and generates electricity; (10) sources include waterfalls or dams. Matches: A, D, E. Option C adds “magically,” which is unsupported by the passage. Ground Truth Answers. A, D, E Model Prediction. A Reason–Answer Mismatch. The model’s reasoning identifies A, D, E as correct, but its final selection includes only A. Options B and F are unsupported; C introduces an extraneous claim (“magically”) not stated in the text.
Overthinking is not helping Long reasoning chains often lead to lower accuracy across GPT-OSS model family. When reasoning token is below 403, GPT OSS 120B achieve 0.65 Exact Match rate. It drops to 0.22 when the number of reasoning token is over 2.8k.
Appendix R LLM Usage
We used large language models (e.g., ChatGPT) solely as assistive tools for (i) light editing of grammar and wording and table reformatting, and (ii) debugging code when running experiments (e.g., clarifying error messages, suggesting fixes). LLMs did not write any source code used in our experiments and did not generate substantive paper content beyond minor edits. All ideas, analyses, experimental designs, and final text are the authors’ own. The authors reviewed and verified all model-assisted edits and take full responsibility for the contents of this paper.